### v3

# Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models

Abstract
Accents are crucial in human communication as they help us understand others and allow us to communicate intelligibly in a way others understand us. While there has been significant progress in ASR, African-accented ASR has been understudied due to a lack of training datasets which are often expensive to create and demand colossal human labor. Our study aims to address this problem by automating the annotation process and reducing annotation-related expenses through informative uncertainty-based data selection. We propose a new multi-rounds adaptation process that uses epistemic uncertainty and evaluate it across several domains, datasets, and high-performing ASR models. Our results show that our approach leads to a 69.44% WER improvement while requiring on average 45% less data than established baselines. Our approach also improves out-of-distribution generalization for very low-resource accents, demonstrating its viability for building generalizable ASR models in the context of accented African ASR. Moreover, the results of our active learning experiments, simulating real-world settings, where there are no gold transcriptions available, also demonstrate the ability of our approach to favor good quality real-life transcriptions. This indicates that our proposed approach addresses the immediate issue of African-accented ASR and has broader implications for improving ASR systems for other underrepresented and low-resource languages and accents. We open-source the code here.
Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models
Bonaventure F. P. Dossou1,2,3,4∗, Atnafu Lambebo Tonja5,∗, Chris Chinenye Emezue1,4,6,∗, Tobi Olatunji7,8,∗ Naome A Etori9,∗, Salomey Osei10,∗, Tosin Adewumi11,∗, Sahib Singh12,∗ ∗Masakhane NLP, 1Mila Quebec AI Institute, 2 McGill University, 3 Lelapa AI, 4 Lanfrica 5Instituto Politécnico Nacional, 6Technical University of Munich, 7Intron Health Inc, 8Georgia Institute of Technology 9University of Minnesota - Twin Cities, 10Universidad de Deusto, 11 Luleå University of Technology, 12 Ford Motor Company
1 Introduction
Generally speaking, Accented English is a form of spoken English that exhibits intonations and pronunciations different from standard English. For instance, Nigerian English differs from British English, which differs from American English. In this work, we consider African-accented English as a form of English spoken by African-native speakers.
Automatic Speech Recognition (ASR) is an active area of research Kodish-Wachs et al. (2018); Finley et al. (2018); Zapata and Kirkedal (2015). In the African context, rich in linguistic diversity and complexity, ASR systems have great potential to reduce language barriers in daily communication. In particular, there is a considerable interest in the development of ASR systems that can accurately distinguish African-accented speech, with a focus on creating systems that can accurately transcribe and interpret speech in a variety of accents and dialects (Koenecke et al., 2020; Hirayama et al., 2012; Tadimeti et al., 2022; Ali et al., 2015; Besacier et al., 2014; Yemmene and Besacier, 2019; Gauthier et al., 2016; Reitmaier et al., 2022; Afonja et al., 2021b). However, developing and deploying ASR systems in an African setting is a challenging task Yemmene and Besacier (2019); De Vries et al. (2014) due to the diversity of African languages, with several dialects and variants characterized by low-resource environments and limited data availability Badenhorst and De Wet (2019, 2017); Barnard et al. (2009); Yemmene and Besacier (2019); DiChristofano et al. (2022). Our methodology facilitates the development of robust, generalizable, and cost-efficient African-accented ASR models. The enhanced generalization for out-of-distribution (OOD) cases, particularly for accents with minimal resources, suggests that our approach extends beyond addressing the specific challenges of African-accented ASR. It holds promise for enhancing ASR systems for other marginalized and resource-scarce languages and accents. This advancement could be a crucial step towards establishing more inclusive and globally accessible ASR technologies. In this work, we reduce this gap, through several contributions listed as follows:
using epistemic uncertainty, we propose a novel multi-rounds adaptation process of several high-performing pretrained ASR models,
we evaluate our approach across several speech domains (clinical, general, both), African-accented speech datasets: AfriSpeech-200 Olatunji et al. (2023), SautiDB Afonja et al. (2021b), MedicalSpeech and CommonVoices English Accented Dataset Ardila et al. (2019), and provide domain-specific as well as accent-specific analyses,
we show that our approach improves the WER performance by 69.44% while reducing significantly the required amount of labeled data (by 35-45%),
we show, based on additional active learning experiments, that our approach is also efficient in real-world settings where there are no gold transcriptions.
2 Challenges and Related Works for African-accented ASR
Although the 2000+ African Languages make up about 31% of the 7000+ languages spoken in the world Eberhard et al. (2019), almost all African languages are classified as "low-resource" Tsvetkov (2017). Despite advancements in ASR over the recent years Kodish-Wachs et al. (2018); Blackley et al. (2019), research or applications in the African context are not well investigated in the literature, likely a product of the paucity of data for accented languages, inadequate research funding, data-hungry approaches well-suited to high-resource settings, and most importantly, data-efficient approaches are inadequately explored. As a result of scarce data and resources, it is also challenging to pretrain large multilingual language models for African languages Dossou et al. (2022).
Despite its potential to improve language communication, ASR models suffer from racial bias and poor performance on accented speech Koenecke et al. (2020), particularly African accents. Recognizing heavy accents in speech data is a challenge for ASR systems since it contains high variability in terms of pronunciations Das et al. (2021). In recent years, we have seen a drastic improvement in ASR systems due to more advanced deep learning systems and the availability of data. It is however shocking that racial disparities and poor performances in accented speech still exist for black speakers. Typically, ASR system generalization is hard to achieve due to difficulties in getting good speech transcript data, including for African accents. The different dialects and accents of many languages introduce some level of difficulty in transcribing such data. These variations are much more complicated in the case of African languages, as speech data in general is more difficult to source. For this reason, there is a poor performance of ASR in African languages.
Improving ASR on African accented speech and linguistic diversity is hence necessary to avoid speakers having to change their way of speaking Mengesha et al. (2021) (a situation that will likely make speakers feel excluded, and the social impact is high), while bringing enormous potential to transform several working areas and daily human needs and communication.
3 Datasets and Methodology
In this work, we established several baselines, across several high-performing ASR systems. Our baselines are produced by feeding the pretrained ASR models, the entirety (100%) of available labeled data. To investigate the effectiveness of our approach, we impose data constraints, choosing fewer training examples than available, to not exceed 60-65% of the entire initial dataset. This simulates a more realistic scenario where the entirety of the data might not be available or feasible to use, thus testing the robustness and efficiency of our method under constrained conditions. We utilize 30% of the initial dataset as our training set and set the remaining 70% as the pool dataset. The purpose of the pool dataset in our study is to allow the model, in our proposed framework, to learn to carefully select the most informative data samples.
To further solidify and demonstrate the effectiveness of our method, we leveraged it in an active learning (AL) setup. AL is a semi-supervised machine learning algorithm that makes use of only a few initial training data points to achieve better performance of a given model. It is important to note that due to the lack of African-accented ASR benchmarks, we primarily compared the results, to our baselines and our best-performing selection technique. Within the defined constraints, for our experiments, for our proposed method, we adjust the sizes of the training and pool datasets, based on the number of training examples available for each domain (clinical, general, and both) (see Tables 2, 4 and Appendix A).
3.1 Datasets
In this work, we primarily used the AfriSpeech-200 dataset Olatunji et al. (2023), a novel 200-hour African-accented English speech corpus carefully curated for clinical and general domains ASR. The dataset covers 120+ African accents, all derived from 5 language families Eberhard et al. (2019): Afro-Asiatic, Indo-European, Khoe-Kwadi (Hainum), Niger-Congo, and Nilo-Saharan, representing regional diversity across western, eastern, and southern Africa. AfriSpeech-200 is a crowd-sourced dataset involving over 2000 African speakers, representing 13 anglophone countries across sub-Saharan Africa and the US (see Table 1).
To establish the dataset-agnostic nature of our approach, we explore three additional datasets:
1.
SautiDB Afonja et al. (2021a): a dataset of Nigerian accent recordings with 919 audio samples, a sampling rate of 48kHz each, for a total amount of 59 minutes of recordings,
2.
MedicalSpeech111https://www.kaggle.com/datasets/paultimothymooney/medical-speech-transcription-and-intent: a dataset which contains 6,661 audio utterances for common medical symptoms like knee pain or headache, for a total of 8 hours of recordings;
3.
CommonVoices English Accented Dataset: a subset of English CommonVoices (version 10) Ardila et al. (2019), with western accents removed as our primary interest is in low-resource settings.
| AfriSpeech Dataset Statistics | |
| Total duration | 200.91 hrs |
| Total clips | 67,577 |
| Unique Speakers | 2,463 |
| Average Audio duration | 10.7 seconds |
| Speaker Gender Ratios - # Clip % | |
| Female | 57.11% |
| Male | 42.41% |
| Other/Unknown | 0.48% |
| Speaker Age Groups - # Clips | |
| <18yrs | 1,264 (1.88%) |
| 19-25 | 36,728 (54.58%) |
| 26-40 | 18,366 (27.29%) |
| 41-55 | 10,374 (15.42%) |
| >56yrs | 563 (0.84%) |
| Clip Domain - # Clips | |
| Clinical | 41,765 (61.80%) |
| General | 25,812 (38.20%) |
| AfriSpeech-200 Dataset Splits | ||||
| Item | Train (*AL Train) | Dev | Test | AL Top-k |
| # Speakers | 1466 | 247 | 750 | ✗ |
| # Hours | 173.4 | 8.74 | 18.77 | ✗ |
| # Accents | 71 | 45 | 108 | ✗ |
| Avg secs/speaker | 425.81 | 127.32 | 90.08 | ✗ |
| clips/speaker | 39.56 | 13.08 | 8.46 | ✗ |
| speakers/accent | 20.65 | 5.49 | 6.94 | ✗ |
| secs/accent | 8791.96 | 698.82 | 625.55 | ✗ |
| # general domain | 21682 (*6504) | 1407 | 2723 | 2000 |
| # clinical domain | 36318 (*10895) | 1824 | 3623 | 3500 |
| # both domain | 58000 (*17400) | 3221 | 6346 | 6500 |
3.2 Methodology
Our methodology is based on epistemic uncertainty. Epistemic Uncertainty (EU) is the uncertainty of the model, most often due to a lack of training data. For a given input , EU is usually defined with the variance (for a given model with some parameters ) as follows:
where . is the learned approximation of . In our experiments, we decided to use the standard deviation defined (STD) as . EU is useful for the model’s training as it improves its robustness and encourages exploration and diversity, to mitigate inductive bias related to underrepresented accents (accents with fewer speech data and speakers).
One of the most popular ways to estimate uncertainty is by inferring predictive distributions with Bayesian Neural Networks (BNNs). A predictive distribution is defined as where is the target, is the input and is the dataset. One can thus inspect the variance (or the standard deviation, as we used in our experiments) and uncover uncertainty, using the predictive distribution. A way of learning the predictive distribution is to learn a distribution over the parameters i.e. the parametric posterior distribution .
There are many ways of computing epistemic uncertainty, but the two most common approaches are Monte Carlo Dropout Gal and Ghahramani (2016) and Deep Ensembles Lakshminarayanan et al. (2017). Many empirical results demonstrated that Deep Ensembles generally work better while being very computationally expensive. Due to computational resource constraints, we decided to demonstrate the effectiveness of our approach using Mc-Dropout empirically. Mc-Dropout provides a scalable way to learn a predictive distribution. It randomly switches off neurons acting as regularizers in a given neural network. Each dropout configuration corresponds to a different sample from the approximate parametric posterior distribution . Sampling from the (approximate) posterior enables Monte Carlo integration of the likelihood of the model, which uncovers the predictive distribution. Our uncertainty-based selection pipeline is detailed in Figure 1 and the algorithm 1. In each adaptation round, we use a finetuned model, to select the most uncertain samples from the pool dataset and add them to the training set for the next round of the model adaptation.
While performing AL experiments, we consider (mask) the samples from the pool dataset to be entirely unlabeled (i.e. no gold transcriptions as targets). This implies that we simulate an unsupervised context, and learn to select the best possible predicted transcription as the target (gold) transcriptions.
Using Mc-Dropout sets a pre-requisite for the model to be equipped with Dropout components during pretraining. This condition eliminates some high-level ASR models like Whisper whisper. However, we finetuned Whisper in our baseline setting and evaluated it against our best-performing approach. This is to assess its robustness and adequacy. In this work, we use the following four state-of-the-art (SOTA) ASR models: Wav2Vec2-XLSR-53 Conneau et al. (2020), Hubert-Large Hsu et al. (2021), WavLM-Large Chen et al. (2022), and NVIDIA Conformer-CTC Large (en-US) Gulati et al. (2020), which we hereafter refer to as Wav2Vec, Hubert, WavLM, and Nemo respectively.
3.3 Experimental Design
Added to the baselines described above, in our framework, we explored two selection techniques: most and random. The most technique consists of selecting the most uncertain audio samples from the pool dataset, while the random one, consists of selecting random audio samples across the pool dataset. In this work, we used the words "sampling", "techniques", and "selection" interchangeably with the same meaning.
We performed various experiments to evaluate our proposed approach’s efficiency and applicability. Each experiment at the model level is performed in three domains: general, clinical, and both (which is the combination of general and clinical data).
For each domain, we have four setups: (a) baseline which consists of simply finetuning a pretrained model on the entire dataset (as described in section 3.1), (b) EU-Most where the pretrained model is finetuned within our framework using the most selection, (c) Random uses the random selection, and (d) AL (w/ EU-Most) where the pretrained model is fine-tuned using active learning and the most technique.
Our experiments are designed to allow us to answer the following specific questions:
1.
given a set of African accents , how does the pretrained ASR model adapt to across the adaptation rounds and domains (clinical, general and both)?
2.
which selection strategy between most and random works better, and for which domain(s)?
3.
which domain(s) help(s) the model to do better, and how does the model perform (in terms of uncertainty) across the domain(s)?
4.
is epistemic uncertainty-based selection effective in low-resource data scenarios?
5.
is uncertainty-based selection, model, and dataset agnostic?
To answer questions 1-4, we define a metric called Uncertainty WER (U-WER). U-WER is described as follows:
where is the speech input, the gold target (transcript), the finetuned ASR model. Due to the nature of the MC-Dropout technique, running it times will produce different speech representations for the original speech input . Each speech representation is denoted , which could consequently lead to different predicted transcription, that we denote by . is the WER function, computing the WER between the predicted transcript and the gold (true) transcript . is the average of WERs between the different speech representations and the target prediction. In our experiments, . The ideal case would be when U-WER0, because this would mean that the model is robust enough to produce consistent transcripts and speech representation instances .
The rationale behind the U-WER is the following: as we acquire new data points that are intrinsically beneficial to the model and its learning, the U-WER should (consistently) decrease or be constant across the multiple adaptation rounds. We could therefore affirm that the model is gaining in robustness, knowledge, and performance; which is a very important aspect of generalization. This last point will be enforced by answering question 5 for which we used the standard WER.
To answer question 5, we evaluated our approach with three additional ASR pretrained models (Nemo, WavLM, and Hubert), and across 3 external datasets (SautiDB, Common Voice English Accented, and MedicalSpeech). Additionally, to answer questions 1-4, for consistency and better visualization, we consider exclusively the top-10 (in terms of frequency) accents present across three adaptation rounds and both selection methods. Regarding the very low-resource settings, we consider the 5 accents with the least number of hours of recordings.
In our AL setup, the U-WER is computed using pairwise WER scores among the predicted transcriptions, rather than using the gold transcriptions. To select the ideal or best predicted transcript for the unlabeled speech , we proceed as described in Algorithm 2.
This implies that for the current speech input , is computed by taking the standard deviation ‘wer_list‘ variable (this corresponds exactly to the formula of the described above). Finally, we assign to , the predicted transcription with the lowest average WER i.e. the lowest value of the dictionary ‘wer_target_dict‘. For our experiments, we use various types of GPUs: 6 RTX 800 (with 48 and 128 GB), and 4 A100 (with 48 and 128 GB). Our training and evaluation were constantly deployed over a month. Our models have 311281844 trainable parameters. Each audio sample has been normalized and processed with a sample rate of 16kHz.
| Model | General | Clinical | Both | |||||||||
| Baseline | Random | EU-Most | AL (w/ EU-Most) | Baseline | Random | EU-Most | AL (w/ EU-Most) | Baseline | Random | EU-Most | AL (w/ EU-Most) | |
| Wav2vec | 0.4980 | 0.1111 | 0.1011 | 0.1059 | 0.5610 | 0.3571 | 0.2457 | 0.2545 | 0.5300 | 0.1666 | 0.1266 | 0.1309 |
| **Hubert | 0.1743 | ✗ | 0.1901 | 0.1887 | 0.2907 | ✗ | 0.2594 | 0.2709 | 0.2365 | ✗ | 0.2453 | 0.2586 |
| **WavLM | 0.1635 | ✗ | 0.1576 | 0.1764 | 0.3076 | ✗ | 0.2313 | 0.2537 | 0.2047 | ✗ | 0.1897 | 0.1976 |
| *Nemo | 0.2824 | ✗ | 0.1765 | 0.1815 | 0.2600 | ✗ | 0.2492 | 0.2526 | 0.3765 | ✗ | 0.2576 | 0.2610 |
| Average Performance | 0.2796 | - | 0.1563 | 0.1631 | 0.3548 | - | 0.2464 | 0.2579 | 0.3369 | - | 0.2043 | 0.2120 |
| Whisper-Medium Olatunji et al. (2023) | 0.2806 | - | - | - | 0.3443 | - | - | - | 0.3116 | - | - | - |
| Dataset | Split and Size for our approach | Finetuning Epochs | Baseline | Uncertainty Sampling w/ most | |||
| Train | Aug | Top-k | Test | (Entire training dataset) | (Train + Aug) | ||
| SautiDB Afonja et al. (2021a) | 234 | 547 | 92 | 138 | 50 | 0.50 | 0.12 |
| MedicalSpeech | 1598 | 3730 | 1333 | 622 | 5 | 0.30 | 0.28 |
| CommonVoices English Accented Dataset (v10.0) Ardila et al. (2019) | 26614 | 62100 | 10350 | 232 | 5 | 0.50 | 0.22 |
| Average | ✗ | ✗ | ✗ | ✗ | ✗ | 0.43 | 0.20 |
4 Results and Discussion
To compute the performance improvement for each domain, we take the ratio of the average of our best method and the baseline. Table 3 presents the results of our experiments. We observe that our uncertainty-based selection approach yields considerable performance gains over the baselines, across all models, all domains, and datasets: general (55.90%), clinical (69.44), and both (61.74%). Additionally, our approach significantly outperforms Whisper-Medium; affirming the effectiveness and reliability of our method. This shows the importance and usefulness of epistemic uncertainty in ASR for low-resource languages.
More importantly, we observe that the most selection provides the best results across all domains. The predominance of most makes sense because, through continual and progressive exposure to the most uncertain samples, the model updates its parameters to become more robust to the noise from uncertain samples. Therefore, it learns to represent those uncertain samples better, yielding a performance improvement. Nevertheless, we speculate that gaps in performance between the general and clinical domains are because clinical samples are the most difficult to capture due to various factors (e.g., domain-specific clinical jargon). This result and analysis provide an answer to question 2, establishing most technique as the most beneficial selection strategy across all domains (see also Figures 2, 3 and 4).
Given the diversity of traits (accents, speakers, gender, age) across the dataset, we attempt to characterize the audio samples and domains that provide the best learning signal to the model. To ground our analysis, we evaluated accent information for top-k uncertain samples. The results are presented in Figures 2, 3 and 4. For further analyses, we focus only on the results of the most technique. For each domain, we find that the top-10 accents are consistent across all rounds (see Figures 2 (a), 3 (a), and 4 (a)), representing the accents with the top-10 highest number of recording hours. With each successive most uncertain top-k samples, we observe that those accents remain the most uncertain and therefore the most difficult to transcribe. From the country and accent statistics (Tables 6, 7 and 8), we observe that those languages have the highest number of speakers and geographical spread across many countries. This interesting observation can be understood as the linguistic richness factor, which potentially explains the high variance induced in the dataset. This implies that, in general, the accents that benefit the model and its learning are the ones that are linguistically rich and diverse. Generally speaking, the model adapts and continuously improves its performance on those accents across the rounds. It is important to notice that for the clinical and both domains (see Figures 3 (a), and 4 (a)), the accent level analysis also reveals some perturbation and noise in the WER variation, which remains constant or slightly increases from Round 2 to Round 3: this suggests that for a critical and complex domain (like the clinical one), there is, and should be, a trade-off between linguistic diversity, rounds, and the value of k samples to be selected. Ultimately, we can positively answer questions 1 and 3, affirming that the model adapts qualitatively and quantitatively well to the beneficial accents and benefits from all domains. Figures 2 (b), 3 (b), and 4 (b) also provide a positive answer to question 4, as we can see a generally improving (or steady) performance on truly low-resource accents in our training dataset, across domains and rounds. This makes our approach very relevant to the low-resourcefulness situation prevalent in many African accents and languages.
Finally, to demonstrate the agnostic aspect of our approach, we evaluated it with three different pretrained ASR models (Hubert, WavLM, and Nemo), and with three datasets containing accented speech on general and clinical domains, only using most selection method. The results are respectively presented in Tables 3 and 4. The main takeaway is that our uncertainty-based adaptation approach performs better than baselines. This efficiently demonstrates that our approach can be applied to any model architecture and with any dataset.
5 Conclusion
We introduced an epistemic uncertainty-based adaptation process for developing efficient, robust African-accented ASR models that perform well across various models, domains, and datasets, particularly in low-resource settings. Our approach effectively identifies and leverages the most informative data samples, enhancing model robustness across different domains.
6 Limitations
6.1 Trade-off between adaptation rounds and new data points selected
In section 4, the results and our analyses revealed that while our approach provided performance improvements, and benefited from linguistically rich accents, we observed that for complex domains (such as clinical and both), it is important to have a stopping criteria and a trade-off between the number of adaptation rounds and the amount of new data points selected at each round. We believe doing such analyses will provide useful insight to improve our current approach.
6.2 Computational Complexity
Similar to machine learning and deep learning approaches, this work was limited by the availability of and access to computational resources. With more computational resources, we would explore Deep Ensembles Lakshminarayanan et al. (2017) as an alternative to our current approach (Mc-Dropout) to estimate epistemic uncertainty from the model.
Appendix A Hyper-parameters
Table 5 shows the hyper-parameter settings used in this study. The top-k value in the table is changed according to the domain used in each of the experiments. For example, when conducting experiments in the general domain, we set the value of top-k to 2k.
| Hyper-parameters | Values |
| attention dropout | 0.1 |
| hidden dropout | 0.1 |
| layer drop | 0.1 |
| train batch size | 16 |
| val batch size | 8 |
| number of epochs | 5 |
| learning rate | 3e-4 |
| maximum audio length | 260000 |
| maximum label length | 260 |
| minimum transcript length | 10 |
| top_k | 2000, 3500, 6500 |
| domains | general, clinical, all |
| active learning rounds | 3 |
| sampling mode | most, random |
| mc-dropout round | 10 |
Appendix B Country Statistics
Table 6 shows the statistics of the countries across the AfriSpeech-200 dataset.
| Country | Clips | Speakers | Duration (seconds) | Duration (hrs) |
| Nigeria | 45875 | 1979 | 512646.88 | 142.40 |
| Kenya | 8304 | 137 | 75195.43 | 20.89 |
| South Africa | 7870 | 223 | 81688.11 | 22.69 |
| Ghana | 2018 | 37 | 18581.13 | 5.16 |
| Botswana | 1391 | 38 | 14249.01 | 3.96 |
| Uganda | 1092 | 26 | 10420.42 | 2.89 |
| Rwanda | 469 | 9 | 5300.99 | 1.47 |
| United States of America | 219 | 5 | 1900.98 | 0.53 |
| Turkey | 66 | 1 | 664.01 | 0.18 |
| Zimbabwe | 63 | 3 | 635.11 | 0.18 |
| Malawi | 60 | 1 | 554.61 | 0.15 |
| Tanzania | 51 | 2 | 645.51 | 0.18 |
| Lesotho | 7 | 1 | 78.40 | 0.02 |
Appendix C Dataset Accents Stats
Tables 7 and 8 provide a list of AfriSpeech accents along with the number of unique speakers, countries where speakers for each accent are located, duration in seconds for each accent, and their presence in the train, dev, and test splits.
| Accent | Clips | Speakers | Duration(s) | Countries | Splits |
| yoruba | 15407 | 683 | 161587.55 | US,NG | train,test,dev |
| igbo | 8677 | 374 | 93035.79 | US,NG,ZA | train,test,dev |
| swahili | 6320 | 119 | 55932.82 | KE,TZ,ZA,UG | train,test,dev |
| hausa | 5765 | 248 | 70878.67 | NG | train,test,dev |
| ijaw | 2499 | 105 | 33178.9 | NG | train,test,dev |
| afrikaans | 2048 | 33 | 20586.49 | ZA | train,test,dev |
| idoma | 1877 | 72 | 20463.6 | NG | train,test,dev |
| zulu | 1794 | 52 | 18216.97 | ZA,TR,LS | dev,train,test |
| setswana | 1588 | 39 | 16553.22 | BW,ZA | dev,test,train |
| twi | 1566 | 22 | 14340.12 | GH | test,train,dev |
| isizulu | 1048 | 48 | 10376.09 | ZA | test,train,dev |
| igala | 919 | 31 | 9854.72 | NG | train,test |
| izon | 838 | 47 | 9602.53 | NG | train,dev,test |
| kiswahili | 827 | 6 | 8988.26 | KE | train,test |
| ebira | 757 | 42 | 7752.94 | NG | train,test,dev |
| luganda | 722 | 22 | 6768.19 | UG,BW,KE | test,dev,train |
| urhobo | 646 | 32 | 6685.12 | NG | train,dev,test |
| nembe | 578 | 16 | 6644.72 | NG | train,test,dev |
| ibibio | 570 | 39 | 6489.29 | NG | train,test,dev |
| pidgin | 514 | 20 | 5871.57 | NG | test,train,dev |
| luhya | 508 | 4 | 4497.02 | KE | train,test |
| kinyarwanda | 469 | 9 | 5300.99 | RW | train,test,dev |
| xhosa | 392 | 12 | 4604.84 | ZA | train,dev,test |
| tswana | 387 | 18 | 4148.58 | ZA,BW | train,test,dev |
| esan | 380 | 13 | 4162.63 | NG | train,test,dev |
| alago | 363 | 8 | 3902.09 | NG | train,test |
| tshivenda | 353 | 5 | 3264.77 | ZA | test,train |
| fulani | 312 | 18 | 5084.32 | NG | test,train |
| isoko | 298 | 16 | 4236.88 | NG | train,test,dev |
| akan (fante) | 295 | 9 | 2848.54 | GH | train,dev,test |
| ikwere | 293 | 14 | 3480.43 | NG | test,train,dev |
| sepedi | 275 | 10 | 2751.68 | ZA | dev,test,train |
| efik | 269 | 11 | 2559.32 | NG | test,train,dev |
| edo | 237 | 12 | 1842.32 | NG | train,test,dev |
| luo | 234 | 4 | 2052.25 | UG,KE | test,train,dev |
| kikuyu | 229 | 4 | 1949.62 | KE | train,test,dev |
| bekwarra | 218 | 3 | 2000.46 | NG | train,test |
| isixhosa | 210 | 9 | 2100.28 | ZA | train,dev,test |
| hausa/fulani | 202 | 3 | 2213.53 | NG | test,train |
| epie | 202 | 6 | 2320.21 | NG | train,test |
| isindebele | 198 | 2 | 1759.49 | ZA | train,test |
| venda and xitsonga | 188 | 2 | 2603.75 | ZA | train,test |
| sotho | 182 | 4 | 2082.21 | ZA | dev,test,train |
| akan | 157 | 6 | 1392.47 | GH | test,train |
| nupe | 156 | 9 | 1608.24 | NG | dev,train,test |
| anaang | 153 | 8 | 1532.56 | NG | test,dev |
| english | 151 | 11 | 2445.98 | NG | dev,test |
| afemai | 142 | 2 | 1877.04 | NG | train,test |
| shona | 138 | 8 | 1419.98 | ZA,ZW | test,train,dev |
| eggon | 137 | 5 | 1833.77 | NG | test |
| luganda and kiswahili | 134 | 1 | 1356.93 | UG | train |
| ukwuani | 133 | 7 | 1269.02 | NG | test |
| sesotho | 132 | 10 | 1397.16 | ZA | train,dev,test |
| benin | 124 | 4 | 1457.48 | NG | train,test |
| kagoma | 123 | 1 | 1781.04 | NG | train |
| nasarawa eggon | 120 | 1 | 1039.99 | NG | train |
| tiv | 120 | 14 | 1084.52 | NG | train,test,dev |
| south african english | 119 | 2 | 1643.82 | ZA | train,test |
| borana | 112 | 1 | 1090.71 | KE | train |
| Accent | Clips | Speakers | Duration(s) | Countries | Splits |
| swahili ,luganda ,arabic | 109 | 1 | 929.46 | UG | train |
| ogoni | 109 | 4 | 1629.7 | NG | train,test |
| mada | 109 | 2 | 1786.26 | NG | test |
| bette | 106 | 4 | 930.16 | NG | train,test |
| berom | 105 | 4 | 1272.99 | NG | dev,test |
| bini | 104 | 4 | 1499.75 | NG | test |
| ngas | 102 | 3 | 1234.16 | NG | train,test |
| etsako | 101 | 4 | 1074.53 | NG | train,test |
| okrika | 100 | 3 | 1887.47 | NG | train,test |
| venda | 99 | 2 | 938.14 | ZA | train,test |
| siswati | 96 | 5 | 1367.45 | ZA | dev,train,test |
| damara | 92 | 1 | 674.43 | NG | train |
| yoruba, hausa | 89 | 5 | 928.98 | NG | test |
| southern sotho | 89 | 1 | 889.73 | ZA | train |
| kanuri | 86 | 7 | 1936.78 | NG | test,dev |
| itsekiri | 82 | 3 | 778.47 | NG | test,dev |
| ekpeye | 80 | 2 | 922.88 | NG | test |
| mwaghavul | 78 | 2 | 738.02 | NG | test |
| bajju | 72 | 2 | 758.16 | NG | test |
| luo, swahili | 71 | 1 | 616.57 | KE | train |
| dholuo | 70 | 1 | 669.07 | KE | train |
| ekene | 68 | 1 | 839.31 | NG | test |
| jaba | 65 | 2 | 540.66 | NG | test |
| ika | 65 | 4 | 576.56 | NG | test,dev |
| angas | 65 | 1 | 589.99 | NG | test |
| ateso | 63 | 1 | 624.28 | UG | train |
| brass | 62 | 2 | 900.04 | NG | test |
| ikulu | 61 | 1 | 313.2 | NG | test |
| eleme | 60 | 2 | 1207.92 | NG | test |
| chichewa | 60 | 1 | 554.61 | MW | train |
| oklo | 58 | 1 | 871.37 | NG | test |
| meru | 58 | 2 | 865.07 | KE | train,test |
| agatu | 55 | 1 | 369.11 | NG | test |
| okirika | 54 | 1 | 792.65 | NG | test |
| igarra | 54 | 1 | 562.12 | NG | test |
| ijaw(nembe) | 54 | 2 | 537.56 | NG | test |
| khana | 51 | 2 | 497.42 | NG | test |
| ogbia | 51 | 4 | 461.15 | NG | test,dev |
| gbagyi | 51 | 4 | 693.43 | NG | test |
| portuguese | 50 | 1 | 525.02 | ZA | train |
| delta | 49 | 2 | 425.76 | NG | test |
| bassa | 49 | 1 | 646.13 | NG | test |
| etche | 49 | 1 | 637.48 | NG | test |
| kubi | 46 | 1 | 495.21 | NG | test |
| jukun | 44 | 2 | 362.12 | NG | test |
| igbo and yoruba | 43 | 2 | 466.98 | NG | test |
| urobo | 43 | 3 | 573.14 | NG | test |
| kalabari | 42 | 5 | 305.49 | NG | test |
| ibani | 42 | 1 | 322.34 | NG | test |
| obolo | 37 | 1 | 204.79 | NG | test |
| idah | 34 | 1 | 533.5 | NG | test |
| bassa-nge/nupe | 31 | 3 | 267.42 | NG | test,dev |
| yala mbembe | 29 | 1 | 237.27 | NG | test |
| eket | 28 | 1 | 238.85 | NG | test |
| afo | 26 | 1 | 171.15 | NG | test |
| ebiobo | 25 | 1 | 226.27 | NG | test |
| nyandang | 25 | 1 | 230.41 | NG | test |
| ishan | 23 | 1 | 194.12 | NG | test |
| bagi | 20 | 1 | 284.54 | NG | test |
| estako | 20 | 1 | 480.78 | NG | test |
| gerawa | 13 | 1 | 342.15 | NG | test |
Appendix D Most common accent distribution
Figures 5 and 6 show the most common accent distribution across the general domain with both random and most selection methods.
Appendix E Ascending and Descending Accents
Figure 7 shows ascending and descending accents across the Top 2k most uncertain samples.

### v4

@@
 Abstract
-Accents are crucial in human communication as they help us understand others and allow us to communicate intelligibly in a way others understand us.
-While there has been significant progress in ASR, African-accented ASR has been understudied due to a lack of training datasets which are often expensive to create and demand colossal human labor.
-Our study aims to address this problem by automating the annotation process and reducing annotation-related expenses through informative uncertainty-based data selection.
-We propose a new multi-rounds adaptation process that uses epistemic uncertainty and evaluate it across several domains, datasets, and high-performing ASR models.
+Accents play a pivotal role in shaping human communication, enhancing our ability to convey and comprehend messages with clarity and cultural nuance.
+While there has been significant progress in Automatic Speech Recognition (ASR), African-accented English ASR has been understudied due to a lack of training datasets, which are often expensive to create and demand colossal human labor.
+Combining several active learning paradigms and the core-set approach, we propose a new multi-rounds adaptation process that uses epistemic uncertainty to automate the annotation process, significantly reducing the associated costs and human labor.
+This novel method streamlines data annotation and strategically selects data samples that contribute most to model uncertainty, thereby enhancing training efficiency.
+We define a new metric called U-WER to track model adaptation to hard accents.
+We evaluate our approach across several domains, datasets, and high-performing speech models.
 Our results show that our approach leads to a 69.44% WER improvement while requiring on average 45% less data than established baselines.
 Our approach also improves out-of-distribution generalization for very low-resource accents, demonstrating its viability for building generalizable ASR models in the context of accented African ASR.
-Moreover, the results of our active learning experiments, simulating real-world settings, where there are no gold transcriptions available, also demonstrate the ability of our approach to favor good quality real-life transcriptions.
-This indicates that our proposed approach addresses the immediate issue of African-accented ASR and has broader implications for improving ASR systems for other underrepresented and low-resource languages and accents.
 We open-source the code here.
-Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models
-Bonaventure F.
-P.
-Dossou1,2,3,4∗, Atnafu Lambebo Tonja5,∗, Chris Chinenye Emezue1,4,6,∗, Tobi Olatunji7,8,∗ Naome A Etori9,∗, Salomey Osei10,∗, Tosin Adewumi11,∗, Sahib Singh12,∗ ∗Masakhane NLP, 1Mila Quebec AI Institute, 2 McGill University, 3 Lelapa AI, 4 Lanfrica 5Instituto Politécnico Nacional, 6Technical University of Munich, 7Intron Health Inc, 8Georgia Institute of Technology 9University of Minnesota - Twin Cities, 10Universidad de Deusto, 11 Luleå University of Technology, 12 Ford Motor Company
 1 Introduction
-Generally speaking, Accented English is a form of spoken English that exhibits intonations and pronunciations different from standard English.
-For instance, Nigerian English differs from British English, which differs from American English.
-In this work, we consider African-accented English as a form of English spoken by African-native speakers.
-Automatic Speech Recognition (ASR) is an active area of research Kodish-Wachs et al.
-(2018); Finley et al.
-(2018); Zapata and Kirkedal (2015).
-In the African context, rich in linguistic diversity and complexity, ASR systems have great potential to reduce language barriers in daily communication.
-In particular, there is a considerable interest in the development of ASR systems that can accurately distinguish African-accented speech, with a focus on creating systems that can accurately transcribe and interpret speech in a variety of accents and dialects (Koenecke et al., 2020; Hirayama et al., 2012; Tadimeti et al., 2022; Ali et al., 2015; Besacier et al., 2014; Yemmene and Besacier, 2019; Gauthier et al., 2016; Reitmaier et al., 2022; Afonja et al., 2021b).
-However, developing and deploying ASR systems in an African setting is a challenging task Yemmene and Besacier (2019); De Vries et al.
-(2014) due to the diversity of African languages, with several dialects and variants characterized by low-resource environments and limited data availability Badenhorst and De Wet (2019, 2017); Barnard et al.
-(2009); Yemmene and Besacier (2019); DiChristofano et al.
-(2022).
-Our methodology facilitates the development of robust, generalizable, and cost-efficient African-accented ASR models.
-The enhanced generalization for out-of-distribution (OOD) cases, particularly for accents with minimal resources, suggests that our approach extends beyond addressing the specific challenges of African-accented ASR.
-It holds promise for enhancing ASR systems for other marginalized and resource-scarce languages and accents.
-This advancement could be a crucial step towards establishing more inclusive and globally accessible ASR technologies.
-In this work, we reduce this gap, through several contributions listed as follows:
-using epistemic uncertainty, we propose a novel multi-rounds adaptation process of several high-performing pretrained ASR models,
-we evaluate our approach across several speech domains (clinical, general, both), African-accented speech datasets: AfriSpeech-200 Olatunji et al.
-(2023), SautiDB Afonja et al.
-(2021b), MedicalSpeech and CommonVoices English Accented Dataset Ardila et al.
-(2019), and provide domain-specific as well as accent-specific analyses,
-we show that our approach improves the WER performance by 69.44% while reducing significantly the required amount of labeled data (by 35-45%),
-we show, based on additional active learning experiments, that our approach is also efficient in real-world settings where there are no gold transcriptions.
-2 Challenges and Related Works for African-accented ASR
-Although the 2000+ African Languages make up about 31% of the 7000+ languages spoken in the world Eberhard et al.
-(2019), almost all African languages are classified as "low-resource" Tsvetkov (2017).
-Despite advancements in ASR over the recent years Kodish-Wachs et al.
-(2018); Blackley et al.
-(2019), research or applications in the African context are not well investigated in the literature, likely a product of the paucity of data for accented languages, inadequate research funding, data-hungry approaches well-suited to high-resource settings, and most importantly, data-efficient approaches are inadequately explored.
-As a result of scarce data and resources, it is also challenging to pretrain large multilingual language models for African languages Dossou et al.
-(2022).
-Despite its potential to improve language communication, ASR models suffer from racial bias and poor performance on accented speech Koenecke et al.
-(2020), particularly African accents.
-Recognizing heavy accents in speech data is a challenge for ASR systems since it contains high variability in terms of pronunciations Das et al.
-(2021).
-In recent years, we have seen a drastic improvement in ASR systems due to more advanced deep learning systems and the availability of data.
-It is however shocking that racial disparities and poor performances in accented speech still exist for black speakers.
-Typically, ASR system generalization is hard to achieve due to difficulties in getting good speech transcript data, including for African accents.
-The different dialects and accents of many languages introduce some level of difficulty in transcribing such data.
-These variations are much more complicated in the case of African languages, as speech data in general is more difficult to source.
-For this reason, there is a poor performance of ASR in African languages.
-Improving ASR on African accented speech and linguistic diversity is hence necessary to avoid speakers having to change their way of speaking Mengesha et al.
-(2021) (a situation that will likely make speakers feel excluded, and the social impact is high), while bringing enormous potential to transform several working areas and daily human needs and communication.
+Automatic Speech Recognition (ASR) is an active research area that powers voice assistant systems (VASs) like Siri and Cortana, enhancing daily communication [27, 17, 46].
+Despite this progress, no current VASs include African languages, which account for about 31% of the world languages, and their unique accents [16, 44].
+This gap underscores the need for ASR systems that can handle the linguistic diversity and complexity of African languages, especially in crucial applications like healthcare.
+Due to the lack of representation of these languages and accents in training data, existing ASR systems often perform inadequately, even mispronouncing African names [36].
+To address these challenges, our work focuses on adapting pretrained speech models to better transcribe African-accented English, defined by unique intonations and pronunciations [7, 22].
+We use epistemic uncertainty (EU) [25] to guide the adaptation process by identifying gaps in model knowledge and prioritizing data for the model to learn from next.
+This is particularly beneficial in scenarios where data annotation is costly or time-consuming, as often seen in the African context [5, 4, 6, 45, 12, 14, 13].
+EU also improves robustness and encourages exploration to mitigate inductive bias from underrepresented accents.
+Common approaches to compute EU include Monte Carlo Dropout (Mc-Dropout) [18] and Deep Ensembles [30], with the latter being more effective but computationally expensive.
+Due to resource constraints, we use Mc-Dropout, which requires models to have dropout components during pretraining.
+Active Learning (AL) is a machine learning (ML) technique that leverages epistemic uncertainty (EU) to select the most informative data points from an unlabeled dataset for labeling, thereby improving model performance with fewer training instances.
+Common types of AL include Deep Bayesian Active Learning (DBAL) [19, 23] and Adversarial Active Learning (AAL) [15].
+AAL selects examples likely to be misclassified by the current model, refining it iteratively by challenging it with difficult cases to enhance robustness.
+The core-set approach (CSA) [42] is also related, as it selects a subset of training data to ensure a model trained on this subset performs comparably to one trained on the full dataset, addressing scalability and efficiency.
+Acquisition functions (AFs) in AL select the most informative samples from an unlabeled dataset for labeling.
+Key AFs include uncertainty sampling (US) [31], Bayesian Active Learning by Disagreement (BALD) [19], and BatchBALD [26].
+US targets data points with the highest model uncertainty.
+BALD maximizes the mutual information between model parameters and predictions.
+BatchBALD is an extension of BALD which selects multiple samples simultaneously but may choose redundant points.
+US is the least computationally expensive, ideal for efficient data labeling.
+In this work, we leverage and combine DBAL, AAL, US, and CSA in the following way (in order):
+1.
+we integrate the CSA by leveraging smaller training subsets ( smaller than the full available training sets)
+- 2.
+3.
+we use the estimated EU, and integrate the idea of AAL by using the US acquisition function.
+We evaluate our approach across several domains (general, clinical, general+clinical aka both), several datasets (AfriSpeech-200 [37], SautiDB [2], MedicalSpeech, CommonVoices English Accented Dataset [3]), and several high-performing speech models (Wav2Vec2-XLSR-53 [10], HuBERT-Large [24], WavLM-Large [8], and NVIDIA Conformer-CTC Large (en-US) [20]).
+Our results show a 69.44% WER improvement while requiring on average 45% less data than established baselines.
+We also adapt the standard Word Error Rate (WER) to create a new metric called Uncertainty WER (U-WER) to track model adaptation to African accents.
+The impact of our approach is substantial.
+It develops more robust, generalizable, and cost-efficient African-accented English ASR models and reduces dependency on large labeled datasets, enabling deployment in various real-world scenarios.
+Our results show improved generalization for out-of-distribution (OOD) cases, especially for accents with minimal resources, addressing specific challenges in African-accented ASR.
+Additionally, by focusing on equitable representation in ASR training, our methodology promotes fairness in AI, ensuring technology serves users across diverse linguistic backgrounds without bias [41, 34, 32].
+Our contributions are listed as follows:
+we combine DBAL, AAL, CSA, and EU to propose a novel way to adapt several high-performing pretrained speech models to build efficient African-accented English ASR models,
+we define a new and simple metric called U-WER that allows us to measure and track how the variance of the model, across hard accents, changes over the adaptation process,
+we show that our approach improves the WER performance by 69.44% while significantly reducing the required amount of labeled data (by 45
+we show, based on additional AL experiments, that our approach is also efficient in real-world settings where there are no gold transcriptions.
+2 Background and Related Works
+2.1 Challenges for African-accented ASR
+State-of-the-art (SOTA) ASR technologies, powered by deep learning and neural network architectures like transformers, achieve high accuracy with Standard American English and major European languages.
+However, they often fail with African accents due to high variability in pronunciation and lack of quality speech data [29, 11].
+This results in racial bias, poor performance, and potential social exclusion as speakers might alter their speech to be understood [29, 28, 9, 33].
+Enhancing ASR for African languages is crucial for equitable voice recognition, especially in healthcare, education, and customer service.
+Solutions should focus on diversifying training datasets and developing robust modeling techniques tailored to these languages unique characteristics.
+2.2 Active Learning
+AL aims at reducing the number of training examples to be labeled by automatically processing the unlabeled examples and then selecting the most informative ones concerning a given cost function for a human to label.
+It is particularly effective when labeled data is scarce or expensive, optimizing the learning process by focusing on samples that most improve the model performance and generalization [43, 19].
+Several works have demonstrated its effectiveness and efficiency.
+An AL setup involves an unlabeled dataset , a labeled training set , and a predictive model with likelihood parameterized by .
+The setup assumes the presence of an oracle to provide predictions for all .
+After training, a batch of data is selected from based on its EU.
+In [21], AL was applied to a toy dataset of How May I Help You recordings.
+Confidence scores were estimated for each word and used to compute the overall confidence score for the audio sample.
+This approach achieved competitive results using 27% less data compared to the baseline.
+In [40], the authors estimated confidence scores for each utterance using an online algorithm with the lattice output of a speech recognizer.
+The utterance scores were filtered through an informativeness function to select an optimal subset of training samples, reducing the labeled data needed for a given WER by over 60%.
+[35] experimented with AL for accent adaptation in speech recognition.
+They adapted a source recognizer to the target accent by selecting a small, matched subset of utterances from a large, untranscribed, multi-accented corpus for human transcription.
+They used a cross-entropy-based relevance measure alongside uncertainty-based sampling.
+However, their experiments on Arabic and English accents showed worse performance compared to baselines, while using more hours of recordings.
 3 Datasets and Methodology
-In this work, we established several baselines, across several high-performing ASR systems.
-Our baselines are produced by feeding the pretrained ASR models, the entirety (100%) of available labeled data.
-To investigate the effectiveness of our approach, we impose data constraints, choosing fewer training examples than available, to not exceed 60-65% of the entire initial dataset.
-This simulates a more realistic scenario where the entirety of the data might not be available or feasible to use, thus testing the robustness and efficiency of our method under constrained conditions.
-We utilize 30% of the initial dataset as our training set and set the remaining 70% as the pool dataset.
-The purpose of the pool dataset in our study is to allow the model, in our proposed framework, to learn to carefully select the most informative data samples.
-To further solidify and demonstrate the effectiveness of our method, we leveraged it in an active learning (AL) setup.
-AL is a semi-supervised machine learning algorithm that makes use of only a few initial training data points to achieve better performance of a given model.
-It is important to note that due to the lack of African-accented ASR benchmarks, we primarily compared the results, to our baselines and our best-performing selection technique.
-Within the defined constraints, for our experiments, for our proposed method, we adjust the sizes of the training and pool datasets, based on the number of training examples available for each domain (clinical, general, and both) (see Tables 2, 4 and Appendix A).
 3.1 Datasets
-In this work, we primarily used the AfriSpeech-200 dataset Olatunji et al.
-(2023), a novel 200-hour African-accented English speech corpus carefully curated for clinical and general domains ASR.
-The dataset covers 120+ African accents, all derived from 5 language families Eberhard et al.
-(2019): Afro-Asiatic, Indo-European, Khoe-Kwadi (Hainum), Niger-Congo, and Nilo-Saharan, representing regional diversity across western, eastern, and southern Africa.
-AfriSpeech-200 is a crowd-sourced dataset involving over 2000 African speakers, representing 13 anglophone countries across sub-Saharan Africa and the US (see Table 1).
-To establish the dataset-agnostic nature of our approach, we explore three additional datasets:
-1.
-SautiDB Afonja et al. (2021a): a dataset of Nigerian accent recordings with 919 audio samples, a sampling rate of 48kHz each, for a total amount of 59 minutes of recordings,
-2.
-MedicalSpeech111https://www.kaggle.com/datasets/paultimothymooney/medical-speech-transcription-and-intent: a dataset which contains 6,661 audio utterances for common medical symptoms like knee pain or headache, for a total of 8 hours of recordings;
-3.
-CommonVoices English Accented Dataset: a subset of English CommonVoices (version 10) Ardila et al. (2019), with western accents removed as our primary interest is in low-resource settings.
+We used the AfriSpeech-200 dataset [37], a 200-hour African-accented English speech corpus for clinical and general ASR.
+This dataset includes 120+ African accents from five language families: Afro-Asiatic, Indo-European, Khoe-Kwadi (Hainum), Niger-Congo, and Nilo-Saharan, representing regional diversity across Africa.
+It was crowd-sourced from over 2000 African speakers from 13 anglophone countries in sub-Saharan Africa and the US.
+To demonstrate the dataset-agnostic nature of our approach, we also explored three additional datasets: (1) SautiDB [1], Nigerian accent recordings with 919 audio samples at a 48kHz sampling rate, totaling 59 minutes; (2) MedicalSpeech111https://www.kaggle.com/datasets/paultimothymooney/medical-speech-transcription-and-intent, containing 6,661 audio utterances of common medical symptoms, totaling 8 hours; and (3) CommonVoices English Accented Dataset, a subset of English Common Voice (version 10) [3], excluding western accents to focus on low-resource settings.
 | AfriSpeech Dataset Statistics | |
@@
 | AfriSpeech-200 Dataset Splits | ||||
-| Item | Train (*AL Train) | Dev | Test | AL Top-k |
+| Item | Train ( Train) | Dev | Test | AL Top-k |
 | # Speakers | 1466 | 247 | 750 | ✗ |
@@
 3.2 Methodology
-Our methodology is based on epistemic uncertainty.
-Epistemic Uncertainty (EU) is the uncertainty of the model, most often due to a lack of training data.
-For a given input , EU is usually defined with the variance (for a given model with some parameters ) as follows:
-where .
-is the learned approximation of .
-In our experiments, we decided to use the standard deviation defined (STD) as .
-EU is useful for the model’s training as it improves its robustness and encourages exploration and diversity, to mitigate inductive bias related to underrepresented accents (accents with fewer speech data and speakers).
-One of the most popular ways to estimate uncertainty is by inferring predictive distributions with Bayesian Neural Networks (BNNs).
-A predictive distribution is defined as where is the target, is the input and is the dataset.
-One can thus inspect the variance (or the standard deviation, as we used in our experiments) and uncover uncertainty, using the predictive distribution.
-A way of learning the predictive distribution is to learn a distribution over the parameters i.e.
-the parametric posterior distribution .
-There are many ways of computing epistemic uncertainty, but the two most common approaches are Monte Carlo Dropout Gal and Ghahramani (2016) and Deep Ensembles Lakshminarayanan et al.
-(2017).
-Many empirical results demonstrated that Deep Ensembles generally work better while being very computationally expensive.
-Due to computational resource constraints, we decided to demonstrate the effectiveness of our approach using Mc-Dropout empirically.
-Mc-Dropout provides a scalable way to learn a predictive distribution.
-It randomly switches off neurons acting as regularizers in a given neural network.
-Each dropout configuration corresponds to a different sample from the approximate parametric posterior distribution .
-Sampling from the (approximate) posterior enables Monte Carlo integration of the likelihood of the model, which uncovers the predictive distribution.
-Our uncertainty-based selection pipeline is detailed in Figure 1 and the algorithm 1.
-In each adaptation round, we use a finetuned model, to select the most uncertain samples from the pool dataset and add them to the training set for the next round of the model adaptation.
-While performing AL experiments, we consider (mask) the samples from the pool dataset to be entirely unlabeled (i.e.
-no gold transcriptions as targets).
-This implies that we simulate an unsupervised context, and learn to select the best possible predicted transcription as the target (gold) transcriptions.
-Using Mc-Dropout sets a pre-requisite for the model to be equipped with Dropout components during pretraining.
-This condition eliminates some high-level ASR models like Whisper whisper.
-However, we finetuned Whisper in our baseline setting and evaluated it against our best-performing approach.
-This is to assess its robustness and adequacy.
-In this work, we use the following four state-of-the-art (SOTA) ASR models: Wav2Vec2-XLSR-53 Conneau et al.
-(2020), Hubert-Large Hsu et al.
-(2021), WavLM-Large Chen et al.
-(2022), and NVIDIA Conformer-CTC Large (en-US) Gulati et al.
-(2020), which we hereafter refer to as Wav2Vec, Hubert, WavLM, and Nemo respectively.
+In our approach, to compute EU for a given input , we perform Mc-Dropout to obtain multiple stochastic forward passes through a finetuned ASR model with likelihood where is the weights of .
+Let be a function that computes the WER between the predicted and the target transcripts.
+Let be the number of stochastic forward passes.
+For each pass , we apply dropout, obtain the output transcript, and compute the WER:
+where is a binary mask matrix sampled independently for each pass. EU(()) can then be estimated from the stochastic forward passes as follows:
+The use of Mc-Dropout requires models to have dropout components during training.
+This excludes some models like Whisper [38], which we still finetuned and evaluated as a baseline.
+We use four state-of-the-art pretrained models: Wav2Vec2-XLSR-53, HuBERT-Large, WavLM-Large, and NVIDIA Conformer-CTC Large (en-US), referred to as Wav2Vec, Hubert, WavLM, and Nemo, respectively.
+3.2.1 Uncertainty WER
+To handle diverse accents, we aim to reduce the EU of the models across hard accents after each adaptation round.
+We define a metric called U-WER to track this.
+To compute U-WER() where is a hard accent, we condition EU on :
+where is the audio sample with accent a and
+Ideally, U-WER0.
+The rationale behind U-WER is that as beneficial data points are acquired, U-WER should decrease or remain constant, indicating increased robustness, knowledge, and performance, crucial for generalization.
+During AL, U-WER is computed using pairwise WER scores among predicted transcriptions, not gold transcriptions (see section 3.3).
+To select the best-generated transcript for unlabeled speech , we follow Algorithm 2.
 3.3 Experimental Design
-Added to the baselines described above, in our framework, we explored two selection techniques: most and random.
-The most technique consists of selecting the most uncertain audio samples from the pool dataset, while the random one, consists of selecting random audio samples across the pool dataset.
-In this work, we used the words "sampling", "techniques", and "selection" interchangeably with the same meaning.
-We performed various experiments to evaluate our proposed approach’s efficiency and applicability.
-Each experiment at the model level is performed in three domains: general, clinical, and both (which is the combination of general and clinical data).
-For each domain, we have four setups: (a) baseline which consists of simply finetuning a pretrained model on the entire dataset (as described in section 3.1), (b) EU-Most where the pretrained model is finetuned within our framework using the most selection, (c) Random uses the random selection, and (d) AL (w/ EU-Most) where the pretrained model is fine-tuned using active learning and the most technique.
-Our experiments are designed to allow us to answer the following specific questions:
+To work within our framework, we define the following selection strategies:
+random: Randomly selects audio samples from .
+EU-Most: Selects the most uncertain audio samples from to add to .
+AL-EU-Most: Combines AL with the EU-Most strategy to finetune the pretrained model.
+We also define standard fine-tuning (SFT) as baseline using all available data for finetuning.
+In SFT, is empty.
+While running the defined strategies in our framework, we impose data constraints, not exceeding 60-65% of the initial dataset after all adaptation rounds.
+is 30% of , and is 70% of .
+This simulates realistic scenarios where not all data might be available, testing the approach’s robustness and efficiency under constraints.
+The number of samples in and is based on available training examples for each domain (see Tables 2, 4, and Appendix A.1).
+Our EU-based pipeline is shown in Figure 1 and Algorithm 1.
+In each adaptation round, we use a finetuned model and a selection strategy to choose samples from to add to .
+During AL experiments, we consider samples from as unlabeled: (1) using Mc-Dropout, we obtain different input representations per audio sample to get different transcripts; (2) we then learn to select the best-generated transcription as the target transcription according to Algorithm 2.
+Our experiments aim to answer the following research questions:
 1.
-given a set of African accents , how does the pretrained ASR model adapt to across the adaptation rounds and domains (clinical, general and both)?
+how does the pretrained ASR model adapt to a set of African accents across adaptation rounds and domains?
 2.
-which selection strategy between most and random works better, and for which domain(s)?
+which selection strategy (EU-most or random) works better, and for which domain(s)?
 3.
-which domain(s) help(s) the model to do better, and how does the model perform (in terms of uncertainty) across the domain(s)?
+which domain(s) help the model perform better, and how does the model perform (in terms of uncertainty) across the domain(s)?
 4.
-is epistemic uncertainty-based selection effective in low-resource data scenarios?
+what is the impact of EU-based selection on the model’s efficiency in low-resource data scenarios?
 5.
 is uncertainty-based selection, model, and dataset agnostic?
-To answer questions 1-4, we define a metric called Uncertainty WER (U-WER). U-WER is described as follows:
-where is the speech input, the gold target (transcript), the finetuned ASR model.
-Due to the nature of the MC-Dropout technique, running it times will produce different speech representations for the original speech input .
-Each speech representation is denoted , which could consequently lead to different predicted transcription, that we denote by .
-is the WER function, computing the WER between the predicted transcript and the gold (true) transcript .
-is the average of WERs between the different speech representations and the target prediction.
-In our experiments, .
-The ideal case would be when U-WER0, because this would mean that the model is robust enough to produce consistent transcripts and speech representation instances .
-The rationale behind the U-WER is the following: as we acquire new data points that are intrinsically beneficial to the model and its learning, the U-WER should (consistently) decrease or be constant across the multiple adaptation rounds.
-We could therefore affirm that the model is gaining in robustness, knowledge, and performance; which is a very important aspect of generalization.
-This last point will be enforced by answering question 5 for which we used the standard WER.
-To answer question 5, we evaluated our approach with three additional ASR pretrained models (Nemo, WavLM, and Hubert), and across 3 external datasets (SautiDB, Common Voice English Accented, and MedicalSpeech).
-Additionally, to answer questions 1-4, for consistency and better visualization, we consider exclusively the top-10 (in terms of frequency) accents present across three adaptation rounds and both selection methods.
-Regarding the very low-resource settings, we consider the 5 accents with the least number of hours of recordings.
-In our AL setup, the U-WER is computed using pairwise WER scores among the predicted transcriptions, rather than using the gold transcriptions.
-To select the ideal or best predicted transcript for the unlabeled speech , we proceed as described in Algorithm 2.
-This implies that for the current speech input , is computed by taking the standard deviation ‘wer_list‘ variable (this corresponds exactly to the formula of the described above).
-Finally, we assign to , the predicted transcription with the lowest average WER i.e.
-the lowest value of the dictionary ‘wer_target_dict‘.
-For our experiments, we use various types of GPUs: 6 RTX 800 (with 48 and 128 GB), and 4 A100 (with 48 and 128 GB).
-Our training and evaluation were constantly deployed over a month.
-Our models have 311281844 trainable parameters.
-Each audio sample has been normalized and processed with a sample rate of 16kHz.
-| Model | General | Clinical | Both | |||||||||
-| Baseline | Random | EU-Most | AL (w/ EU-Most) | Baseline | Random | EU-Most | AL (w/ EU-Most) | Baseline | Random | EU-Most | AL (w/ EU-Most) | |
-| Wav2vec | 0.4980 | 0.1111 | 0.1011 | 0.1059 | 0.5610 | 0.3571 | 0.2457 | 0.2545 | 0.5300 | 0.1666 | 0.1266 | 0.1309 |
-| **Hubert | 0.1743 | ✗ | 0.1901 | 0.1887 | 0.2907 | ✗ | 0.2594 | 0.2709 | 0.2365 | ✗ | 0.2453 | 0.2586 |
-| **WavLM | 0.1635 | ✗ | 0.1576 | 0.1764 | 0.3076 | ✗ | 0.2313 | 0.2537 | 0.2047 | ✗ | 0.1897 | 0.1976 |
-| *Nemo | 0.2824 | ✗ | 0.1765 | 0.1815 | 0.2600 | ✗ | 0.2492 | 0.2526 | 0.3765 | ✗ | 0.2576 | 0.2610 |
-| Average Performance | 0.2796 | - | 0.1563 | 0.1631 | 0.3548 | - | 0.2464 | 0.2579 | 0.3369 | - | 0.2043 | 0.2120 |
-| Whisper-Medium Olatunji et al. (2023) | 0.2806 | - | - | - | 0.3443 | - | - | - | 0.3116 | - | - | - |
-| Dataset | Split and Size for our approach | Finetuning Epochs | Baseline | Uncertainty Sampling w/ most | |||
-| Train | Aug | Top-k | Test | (Entire training dataset) | (Train + Aug) | ||
-| SautiDB Afonja et al. (2021a) | 234 | 547 | 92 | 138 | 50 | 0.50 | 0.12 |
-| MedicalSpeech | 1598 | 3730 | 1333 | 622 | 5 | 0.30 | 0.28 |
-| CommonVoices English Accented Dataset (v10.0) Ardila et al. (2019) | 26614 | 62100 | 10350 | 232 | 5 | 0.50 | 0.22 |
-| Average | ✗ | ✗ | ✗ | ✗ | ✗ | 0.43 | 0.20 |
+U-WER will answer question 4.
+To answer question 5, we evaluated our approach with three additional pretrained models (Nemo, WavLM, and Hubert) and across three external datasets (SautiDB, CommonVoices English Accented Dataset, and MedicalSpeech).
+For consistency and better visualization, we considered the top-10 (in terms of frequency) accents across three adaptation rounds and both selection strategies to answer questions 1-4.
+For very low-resource settings, we considered the five accents with the least recording hours.
+We used various GPUs for our experiments: 6 RTX 800, and 4 A100.
+Training and evaluation were conducted over a month.
+Our models have approximately 311 million trainable parameters.
+Each audio sample was normalized and processed at a 16kHz sample rate.
+We used default parameters from the HuggingFace library for each pretrained model.
+| Model | General | Clinical | Both | ||||||
+| Baseline | EU-Most | AL-EU-Most | Baseline | EU-Most | AL-EU-Most | Baseline | EU-Most | AL-EU-Most | |
+| Wav2vec | 0.4980 | 0.1011 | 0.1059 | 0.5610 | 0.2457 | 0.2545 | 0.5300 | 0.1266 | 0.1309 |
+| **Hubert | 0.1743 | 0.1901 | 0.1887 | 0.2907 | 0.2594 | 0.2709 | 0.2365 | 0.2453 | 0.2586 |
+| **WavLM | 0.1635 | 0.1576 | 0.1764 | 0.3076 | 0.2313 | 0.2537 | 0.2047 | 0.1897 | 0.1976 |
+| *Nemo | 0.2824 | 0.1765 | 0.1815 | 0.2600 | 0.2492 | 0.2526 | 0.3765 | 0.2576 | 0.2610 |
+| Average Performance | 0.2796 | 0.1563 | 0.1631 | 0.3548 | 0.2464 | 0.2579 | 0.3369 | 0.2043 | 0.2120 |
+| Whisper-Medium | 0.2806 | - | - | 0.3443 | - | - | 0.3116 | - | - |
 4 Results and Discussion
-To compute the performance improvement for each domain, we take the ratio of the average of our best method and the baseline.
-Table 3 presents the results of our experiments.
-We observe that our uncertainty-based selection approach yields considerable performance gains over the baselines, across all models, all domains, and datasets: general (55.90%), clinical (69.44), and both (61.74%).
-Additionally, our approach significantly outperforms Whisper-Medium; affirming the effectiveness and reliability of our method.
-This shows the importance and usefulness of epistemic uncertainty in ASR for low-resource languages.
-More importantly, we observe that the most selection provides the best results across all domains.
-The predominance of most makes sense because, through continual and progressive exposure to the most uncertain samples, the model updates its parameters to become more robust to the noise from uncertain samples.
-Therefore, it learns to represent those uncertain samples better, yielding a performance improvement.
-Nevertheless, we speculate that gaps in performance between the general and clinical domains are because clinical samples are the most difficult to capture due to various factors (e.g., domain-specific clinical jargon).
-This result and analysis provide an answer to question 2, establishing most technique as the most beneficial selection strategy across all domains (see also Figures 2, 3 and 4).
-Given the diversity of traits (accents, speakers, gender, age) across the dataset, we attempt to characterize the audio samples and domains that provide the best learning signal to the model.
-To ground our analysis, we evaluated accent information for top-k uncertain samples.
-The results are presented in Figures 2, 3 and 4.
-For further analyses, we focus only on the results of the most technique.
-For each domain, we find that the top-10 accents are consistent across all rounds (see Figures 2 (a), 3 (a), and 4 (a)), representing the accents with the top-10 highest number of recording hours.
-With each successive most uncertain top-k samples, we observe that those accents remain the most uncertain and therefore the most difficult to transcribe.
-From the country and accent statistics (Tables 6, 7 and 8), we observe that those languages have the highest number of speakers and geographical spread across many countries.
-This interesting observation can be understood as the linguistic richness factor, which potentially explains the high variance induced in the dataset.
-This implies that, in general, the accents that benefit the model and its learning are the ones that are linguistically rich and diverse.
-Generally speaking, the model adapts and continuously improves its performance on those accents across the rounds.
-It is important to notice that for the clinical and both domains (see Figures 3 (a), and 4 (a)), the accent level analysis also reveals some perturbation and noise in the WER variation, which remains constant or slightly increases from Round 2 to Round 3: this suggests that for a critical and complex domain (like the clinical one), there is, and should be, a trade-off between linguistic diversity, rounds, and the value of k samples to be selected.
-Ultimately, we can positively answer questions 1 and 3, affirming that the model adapts qualitatively and quantitatively well to the beneficial accents and benefits from all domains.
-Figures 2 (b), 3 (b), and 4 (b) also provide a positive answer to question 4, as we can see a generally improving (or steady) performance on truly low-resource accents in our training dataset, across domains and rounds.
-This makes our approach very relevant to the low-resourcefulness situation prevalent in many African accents and languages.
-Finally, to demonstrate the agnostic aspect of our approach, we evaluated it with three different pretrained ASR models (Hubert, WavLM, and Nemo), and with three datasets containing accented speech on general and clinical domains, only using most selection method.
-The results are respectively presented in Tables 3 and 4.
-The main takeaway is that our uncertainty-based adaptation approach performs better than baselines.
-This efficiently demonstrates that our approach can be applied to any model architecture and with any dataset.
+To compute the performance improvement for each domain, we take the ratio of the average of our best strategy and the baseline.
+Table 3 shows the results of our experiments, indicating that our uncertainty-based selection approach significantly outperforms the baselines across all models, domains, and datasets: general (55.90%), clinical (69.44%), and both (61.74%).
+Our approach also surpasses Whisper-Medium [37, 39], demonstrating the importance of epistemic uncertainty in ASR for low-resource languages.
+The EU-Most selection strategy proves to be the most effective across all domains due to the model exposure to highly uncertain samples, enhancing robustness and performance.
+However, performance disparities between general and clinical domains are noted, likely due to clinical samples complexity.
+These findings confirm EU-Most as the superior selection strategy, as detailed in the results and illustrated in Figures 2, 3, and 4.
+This answers question 2.
+To identify the best learning signals within a diverse dataset characterized by various accents, speaker traits, genders, and ages, we analyzed the top-k uncertain accents using the EU-Most selection strategy.
+Our findings, illustrated in Figures 2, 3, and 4, show that the top-10 accents (most represented in recording hours) remained consistently challenging across all rounds of analysis (refer to Figures 2, 3, 4 and Tables 6, 7, and 8).
+These accents, characterized by high linguistic richness and variability, aid in model learning and enhance performance over time.
+We positively answer questions 1 and 3, confirming that the model adapts effectively to the beneficial accents from all domains.
+This demonstrates that the model adapts qualitatively and quantitatively well to the beneficial accents and benefits from all domains.
+Figures 2 (b), 3 (b), and 4 (b) also affirm positive outcomes for question 4, showing consistent improvement or stable performance on low-resource accents.
+This highlights the relevance of our approach in addressing the challenges associated with the low resource availability typical of many African accents and languages.
+To demonstrate the agnostic aspect of our approach, we evaluated it with three additional pretrained models (Hubert, WavLM, and Nemo) and three datasets containing accented speech in general and clinical domains, using only the EU-Most selection strategy.
+The results, shown in Tables 3 and 4, indicate that our uncertainty-based adaptation approach consistently outperforms baselines.
+This confirms that our approach applies to any model architecture and dataset, and allows us to answer positively question 5.
 5 Conclusion
-We introduced an epistemic uncertainty-based adaptation process for developing efficient, robust African-accented ASR models that perform well across various models, domains, and datasets, particularly in low-resource settings.
-Our approach effectively identifies and leverages the most informative data samples, enhancing model robustness across different domains.
+We combined several AL paradigms, the CSA, and EU to create a novel multi-round adaptation process for high-performing pretrained speech models, aiming to build efficient African-accented English ASR models.
+We introduced a new metric called U-WER to track model adaptation to hard accents.
+Our experiments showed a remarkable 69.44% WER improvement while reducing the data required for effective training by approximately 45% compared to existing baselines.
+This demonstrates our approach’s efficiency and potential to significantly lower barriers to ASR technologies in underserved regions.
+Our method enhances model robustness and generalization across various domains, datasets, and accents, crucial for scalable ASR systems.
+This also helps mitigate bias in ASR technologies, promoting more inclusive and fair AI applications.
 6 Limitations
-6.1 Trade-off between adaptation rounds and new data points selected
-In section 4, the results and our analyses revealed that while our approach provided performance improvements, and benefited from linguistically rich accents, we observed that for complex domains (such as clinical and both), it is important to have a stopping criteria and a trade-off between the number of adaptation rounds and the amount of new data points selected at each round.
-We believe doing such analyses will provide useful insight to improve our current approach.
-6.2 Computational Complexity
-Similar to machine learning and deep learning approaches, this work was limited by the availability of and access to computational resources.
-With more computational resources, we would explore Deep Ensembles Lakshminarayanan et al.
-(2017) as an alternative to our current approach (Mc-Dropout) to estimate epistemic uncertainty from the model.
-Appendix A Hyper-parameters
+In discussing trade-offs (Section 4), we observed that while our approach improves performance, especially with linguistically rich accents, it is crucial for complex domains like the clinical and both domains to establish a stopping criterion that balances adaptation rounds with the selection of new data points each round.
+This balance is vital for refining our methodology.
+Regarding computational complexity, our work faced limitations due to the availability of computational resources.
+With better resources, we would consider implementing Deep Ensembles [30] as an alternative to our current Mc-Dropout method for estimating epistemic uncertainty.
+Acknowledgments and Disclosure of Funding
+The author is very grateful to Ines Arous for her help in reviewing the manuscript and suggesting very important changes to improve the quality of the paper.
+The author acknowledges the support of the Mila Quebec AI Institute for computing resources.
+The author acknowledges Intron Health for providing the Afrispeech-200 dataset.
+The author is grateful to Atnafu Lambebo Tonja, Chris Chinenye Emezue, Tobi Olatunji, Naome A Etori, Salomey Osei, Tosin Adewumi, and Sahib Singh; for their help in the early stage of this project.
+Checklist
+1.
+For all authors…
+Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] (See Section 4)
+Did you describe the limitations of your work? [Yes] (See Section 6)
+Did you discuss any potential negative societal impacts of your work? [N/A]
+Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
+2.
+If you are including theoretical results…
+Did you state the full set of assumptions of all theoretical results? [N/A]
+Did you include complete proofs of all theoretical results? [N/A]
+3.
+If you ran experiments (e.g. for benchmarks)…
+Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] (See the abstract)
+- (b)
+- (c)
+Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] (See Section 3.3)
+4.
+If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
+If your work uses existing assets, did you cite the creators? [Yes] (See Section 3.1)
+Did you mention the license of the assets? [Yes] (See Section 3.1)
+Did you include any new assets either in the supplemental material or as a URL? [N/A]
+Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
+Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
+5.
+If you used crowdsourcing or conducted research with human subjects…
+Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
+Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
+Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
+Appendix A Appendices
+A.1 Hyper-parameters
 Table 5 shows the hyper-parameter settings used in this study.
@@
 | active learning rounds | 3 |
-| sampling mode | most, random |
+| sampling mode | EU-Most, random |
 | Mc-dropout round | 10 |
-Appendix B Country Statistics
+A.2 Country Statistics
 Table 6 shows the statistics of the countries across the AfriSpeech-200 dataset.
@@
 | Lesotho | 7 | 1 | 78.40 | 0.02 |
-Appendix C Dataset Accents Stats
+A.3 Dataset Accents Stats
 Tables 7 and 8 provide a list of AfriSpeech accents along with the number of unique speakers, countries where speakers for each accent are located, duration in seconds for each accent, and their presence in the train, dev, and test splits.
@@
 | gerawa | 13 | 1 | 342.15 | NG | test |
-Appendix D Most common accent distribution
-Figures 5 and 6 show the most common accent distribution across the general domain with both random and most selection methods.
-Appendix E Ascending and Descending Accents
+A.4 Most common accent distribution
+Figures 5 and 6 show the most common accent distribution across the general domain with both random and EU-Most selection strategies.
+A.5 Ascending and Descending Accents
 Figure 7 shows ascending and descending accents across the Top 2k most uncertain samples.